# On the Need to Rethink Trust in AI Assistants for Software Development: A Critical Review

### v2

@@ III-A2 Critical review of recent SE literature
A critical review focuses on evaluating key issues and analyzing a sample of studies with shared characteristics [12]. The issue we intended to evaluate is the lack of trust conceptualization in SE research (related to RQ2). To this end, we used dblp SPARQL to retrieve all articles that met the inclusion/exclusion criteria:
  • Include papers in leading SE conferences (ICSE, ASE, FSE) and journals (TSE, TOSEM) published recently, i.e., just before or after the GenAI hype (2022-2024).
@@ III-A3 Data availability
We provide all retrieval and analysis scripts and all metadata we retrieved from Google Scholar and dblp as part of our supplementary material [13]. Moreover, we share the annotated lists of papers, including exclusion reasons, and all extracted excerpts. All included research papers are referenced in Table II.
@@ III-B Trust Conceptualizations across Disciplines
In this section, we report the results of our literature review across disciplines. Note that with this review, our goal was not to achieve a complete coverage of all articles that describe conceptualizations of trust in the target disciplines. Our goal was to identify whether articles in the selected disciplines commonly refer to established trust models, with the goal of comparing SE articles to related disciplines. Although we acknowledge this as a limitation of our research, the difference between the SE articles and the HCI/IS articles was so clear that our observations would most likely not change considerably when scaling up the analysis. Table I lists the trust models that at least two articles referred to. The models proposed by [95] [95] and [80] [80] are used across HCI and IS, while [100] [100] and [46] [46] are IS-specific. The other models listed in the table are primarily referenced in HCI, but less frequently than the [95] [95] and [80] [80] models.
@@
(1995) [95]
Multi-disciplinary model of organizational trust. [139, 52, 117, 114, 33, 28, 109] [139, 52, 15, 114] [83, 138, 167, 144] [146, 126, 36, 7]
@@
(2004) [80]
Trust in automation systems and technology. [139, 114, 28, 109, 156] [139, 15, 142, 114, 165] [138] [146, 7]
@@
(2011) [98]
Model of trust in technology grounded in system-level traits [139, 52] [139, 52] None [146]
@@
(2019) [52]
Empirically validated model for measuring intelligent systems trust. [139, 114] [139, 114] None None
@@
(1995) [96]
Dual-factor model of organizational interpersonal trust. [109, 117] [15] None None
@@
(1998) [118]
Integration of organizational trust across disciplines. [117, 28] None [32] None
@@
(2000) [90]
Human-computer trust model, focus on affective & cognitive trust. [109, 78] [15] None None
@@
(2002) [100]
Model of initial trust in e-commerce settings. None None [83, 145, 32, 57, 82] [152]
@@
(2003) [46]
Model of trust formation in e-commerce settings. None None [138, 57, 143] [150]
@@
HCI [139, 27, 52, 133, 1, 117, 78, 114, 33, 61, 28, 67, 109] [65, 51, 8, 105] None [156]
HCI-AI [133, 139, 61, 160, 52, 28, 60, 15, 135, 45, 142, 114, 165, 6, 134] [28, 8, 102] [69] [132]
IS [83, 138, 145, 32, 57, 50, 167, 110, 5, 82, 92, 143, 144] None [17, 18] None
IS-AI [122, 146, 126, 36, 54, 35, 150, 76, 64, 148, 7] [71, 88, 49, 163] [89] [152]
SE [23, 63, 55, 66, 85, 21, 29, 72, 155, 149, 20, 16] [59, 26] None None
SE-AI [3, 66] [86, 168, 71, 81, 88] [89] None
SE-CR [108, 4, 158, 77, 22] None None None
@@ III-B1 Information Systems and Management (IS)
In IS research, trust has been studied since the 1990s [137]. The articles we included in our review of IS literature were published in the time span from 2000 to 2024. Of the 24 research papers we analyzed, only four did not embed their studies in existing trust models. It was common for IS papers to customize existing trust models (13 out of 25 articles).
@@
Several articles discussed that trust is a dynamic concept that develops over time [83]. The factors and processes that lead to trust formation change gradually [95], and can include aspects such as people’s perceptions about a company or brand [138]. This means that trust in a company can transfer to trust in a particular tool, a concept called trust transference [141, 153]. [122] applied this concept to AI agents and noted that trust in an AI creator can be transferred to its AI agent when there is a perceived meaningful association between them [122].
@@
IS articles clearly distinguish initial trust, which is important to create a trustworthy first impression of a new system and encourage adoption [83], from the before-mentioned process of trust formation and later stages such as knowledge-based trust, where the individual knows the other party well enough to predict the party’s behavior in a situation [146]. From a tool perspective, the later phases of trust are sometimes referred to as postadoption, where aspects such as predictability and helpfulness of a tool [116] play an important role [145]. [100]’s model is frequently referred to as a common IS model for initial trust, whereas [46]’s model focuses on trust formation.
@@
In [100]’s model, based on [95], trust in vendors of e-commerce websites is a multidimensional construct centered around trusting beliefs, that is, perceptions of the competence, benevolence, and integrity of the vendor and trusting intentions, that is, the willingness to depend or, in other words, the decision to make oneself vulnerable to the vendor [100]. In addition to a trust model, [100] also introduced instruments to measure the concepts in their model (see Figure 1 and Section IV).
@@
[46] models consumers’ online purchase intentions as a combination of perceived usefulness, ease-of-use, and trust in the vendor [46]. They show that online trust is built through (1) a belief that the vendor has nothing to gain by cheating, (2) a belief that there are safety mechanisms built into the web site, and (3) by having a typical interface, (4) one that is, moreover, easy to use.
@@
An important observation by [95] is repeated in IS articles: humans differ in their general tendency to be willing to trust others [83], that is, their propensity [95, 32] or disposition to trust [82]. These individual differences are a potential confounding factor in studies with developers investigating trust in AI assistants.
@@
Two other central constructs are trusting intention, which is the trustor’s willingness to depend on the trustee [83] and a person’s trusting beliefs [146], that is favorable object-specific beliefs [145] that capture the trustor’s perception of whether the trustee has the desired attributes to be trusted [83] (see also [100]’s model). Moreover, IS literature discusses how the trusting beliefs proposed by [95] (benevolence, ability, and integrity) can foster technology acceptance [19]. [150] categorize those trusting beliefs into cognitive trust and affective trust [150] (see also [75] [75]). Cognitive trust captures a customer’s confidence or willingness to rely on a service provider’s competence and reliability while affective trust is the confidence that one places in another party on the basis of feelings generated by the level of care and concern as demonstrated by the other party. In summary, competence (i.e., ability in [95]’s model) and predictability influence cognitive trust while benevolence and integrity influence affective trust [150].
@@
The IS literature discusses the validity of applying models of human trust to technology [83, 32, 122], which is referred to as plausibility of trust in philosophy (see Section II-C). An aspect to consider is that humans have long been known to anthropomorphize technology [106]. To study trust relationships between humans and human-like tools, researchers have adapted the concept of interpersonal trust to technology, for example, to study trust in automation or autonomous systems (see the description of [80]’s trust model in Section III-B2). However, researchers have also argued that such models do not necessarily apply to GenAI tools due to their nondeterminism [7].
@@
IS articles further discuss that increasing automation and complexity leads to systems being perceived as opaque and sophisticated, hence trust becomes more and more important [80]. This situation leads to uncertainty and a perceived risk of using an automation tool, which could hinder adoption [25, 145]. However, [32] also write that trust can only arise when there exists a state of dependence between the trustor and trustee, and when acting on this dependence entails risk, i.e., the trustor possesses uncertainty about the outcomes and vulnerability to a potential loss if the outcomes are undesirable[32]. This is certainly true for software engineers who use GenAI tools to develop software. The role of risk and vulnerability appears also in the psychology literature, in particular [95] (see Figure 1), as well as in the philosophical debate summarized in Section II-B.
@@
In the context of AI for decision-making, both mistakenly denied trust and unfounded trust can occur [126]. [126] summarize these situations as follows: If humans increasingly leverage AI to inform, derive, and justify decisions, it also becomes important to quantify when, how, why, and under which conditions they tend to overly trust or mistrust those systems [126]. This view corresponds to the concepts of undertrust and overtrust as described by [80] [80].
@@
Regarding trust quantification, [126] note that trust is measured differently in different research fields [126]. However, they also mention that a common approach used across disciplines is to ask study subjects about their trusting beliefs or trusting intentions towards a certain entity, for which standardized survey instruments exist (cf. [47]). An exemplary operationalization of trust in decision-making based on observed behavior is calculating the probability that users follow a model’s predictions by utilizing usage data [115, 125]. This resembles our motivating example of software developers accepting code suggestions. We discuss trust operationalization and measurement in more detail in Section IV.

### v3

@@ III-A2 Critical review of recent SE literature
A critical review focuses on evaluating key issues and analyzing a sample of studies with shared characteristics [baltes2022sampling]. The issue we intended to evaluate is the lack of trust conceptualization in SE research (related to RQ2). To this end, we used dblp SPARQL to retrieve all articles that met the inclusion/exclusion criteria:
  • Include papers in leading SE conferences (ICSE, ASE, FSE) and journals (TSE, TOSEM) published recently, i.e., just before or after the GenAI hype (2022–2024).
@@ III-A3 Data availability
We provide all retrieval and analysis scripts and all metadata we retrieved from Google Scholar and dblp as part of our supplementary material [baltes_2025_15130575]. Moreover, we share the annotated lists of papers, including exclusion reasons, and all extracted excerpts. All included research papers are referenced in Table II.
@@ III-B Trust Conceptualizations across Disciplines
In this section, we report the results of our literature review across disciplines. Note that with this review, our goal was not to achieve a complete coverage of all articles that describe conceptualizations of trust in the target disciplines. Our goal was to identify whether articles in the selected disciplines commonly refer to established trust models, with the goal of comparing SE articles to related disciplines. Although we acknowledge this as a limitation of our research, the difference between the SE articles and the HCI/IS articles was so clear that our observations would most likely not change considerably when scaling up the analysis. Table I lists the trust models that at least two articles referred to. The models proposed by [Mayer1995] [Mayer1995] and [Lee2004Trust] [Lee2004Trust] are used across HCI and IS, while [DBLP:journals/jsis/McKnightCK02] [DBLP:journals/jsis/McKnightCK02] and [DBLP:journals/misq/GefenKS03] [DBLP:journals/misq/GefenKS03] are IS-specific. The other models listed in the table are primarily referenced in HCI, but less frequently than the [Mayer1995] [Mayer1995] and [Lee2004Trust] [Lee2004Trust] models.
@@
(1995) [Mayer1995]
Multi-disciplinary model of organizational trust. [DBLP:journals/computer/SousaLCM24, DBLP:journals/behaviourIT/Gulati0L19, DBLP:journals/ijmms/RiegelsbergerSM05, pinto2022trust, DBLP:journals/ijmms/CorritoreKW03a, Cai02012025, DBLP:conf/compsac/NothdurftBLM12] [DBLP:journals/computer/SousaLCM24, DBLP:journals/behaviourIT/Gulati0L19, DBLP:journals/pacmhci/BanovicYRL23, pinto2022trust] [DBLP:journals/jsis/LiHV08, DBLP:journals/ejis/SollnerHL16, Zhang2018, Tansley2007] [DBLP:journals/electronicmarkets/ThiebesLS21, DBLP:journals/jds/SchmidtBT20, Dennis03042023, info:doi/10.2196/15154]
@@
(2004) [Lee2004Trust]
Trust in automation systems and technology. [DBLP:journals/computer/SousaLCM24, pinto2022trust, Cai02012025, DBLP:conf/compsac/NothdurftBLM12, DBLP:conf/chi/WidderDHHD21] [DBLP:journals/computer/SousaLCM24, DBLP:journals/pacmhci/BanovicYRL23, DBLP:journals/chb/SuenH23, pinto2022trust, DBLP:journals/ijhci/YokoiEFN21] [DBLP:journals/ejis/SollnerHL16] [DBLP:journals/electronicmarkets/ThiebesLS21, info:doi/10.2196/15154]
@@
(2011) [DBLP:journals/tmis/McKnightCTC11]
Model of trust in technology grounded in system-level traits. [DBLP:journals/computer/SousaLCM24, DBLP:journals/behaviourIT/Gulati0L19] [DBLP:journals/computer/SousaLCM24, DBLP:journals/behaviourIT/Gulati0L19] None [DBLP:journals/electronicmarkets/ThiebesLS21]
@@
(2019) [DBLP:journals/behaviourIT/Gulati0L19]
Empirically validated model for measuring intelligent systems trust. [DBLP:journals/computer/SousaLCM24, pinto2022trust] [DBLP:journals/computer/SousaLCM24, pinto2022trust] None None
@@
(1995) [mcallister1995affect]
Dual-factor model of organizational interpersonal trust. [DBLP:conf/compsac/NothdurftBLM12, DBLP:journals/ijmms/RiegelsbergerSM05] [DBLP:journals/pacmhci/BanovicYRL23] None None
@@
(1998) [rousseau1998not]
Integration of organizational trust across disciplines. [DBLP:journals/ijmms/RiegelsbergerSM05, Cai02012025] None [DBLP:conf/hicss/ChopraW03] None
@@
(2000) [madsen2000measuring]
Human-computer trust model, focus on affective & cognitive trust. [DBLP:conf/compsac/NothdurftBLM12, DBLP:conf/um/KrausWM20] [DBLP:journals/pacmhci/BanovicYRL23] None None
@@
(2002) [DBLP:journals/jsis/McKnightCK02]
Model of initial trust in e-commerce settings. None None [DBLP:journals/jsis/LiHV08, DBLP:journals/tem/ThatcherMBAR11, DBLP:conf/hicss/ChopraW03, DBLP:journals/jcis/HoehleHG12, DBLP:journals/db/LiHV06] [DBLP:journals/isci/VirvouTT24]
@@
(2003) [DBLP:journals/misq/GefenKS03]
Model of trust formation in e-commerce settings. None None [DBLP:journals/ejis/SollnerHL16, DBLP:journals/jcis/HoehleHG12, Tam2020] [Troshani03092021]
@@
HCI [DBLP:journals/computer/SousaLCM24, BURGOON2000553, DBLP:journals/behaviourIT/Gulati0L19, DBLP:journals/ijhci/Shneiderman20, DBLP:conf/chi/AdarTT13, DBLP:journals/ijmms/RiegelsbergerSM05, DBLP:conf/um/KrausWM20, pinto2022trust, DBLP:journals/ijmms/CorritoreKW03a, DBLP:journals/chb/HuoZYSH22, Cai02012025, DBLP:journals/hhci/JoinsonRBS10, DBLP:conf/compsac/NothdurftBLM12] [DBLP:journals/ijhci/Jeon24, GULATI2024100495, DBLP:journals/ijhci/BachKHBS24, DBLP:journals/csur/MyersHC96] None [DBLP:conf/chi/WidderDHHD21]
HCI-AI [DBLP:journals/ijhci/Shneiderman20, DBLP:journals/computer/SousaLCM24, DBLP:journals/chb/HuoZYSH22, DBLP:journals/ijhci/XuDGG23, DBLP:journals/behaviourIT/Gulati0L19, Cai02012025, DBLP:journals/ijhci/HuoZQYYYS24, DBLP:journals/pacmhci/BanovicYRL23, DBLP:journals/ijhci/SilvaSHGG23, DBLP:journals/ijhci/GaribayWAABCFFG23, DBLP:journals/chb/SuenH23, pinto2022trust, DBLP:journals/ijhci/YokoiEFN21, DBLP:conf/chi/AmershiWVFNCSIB19, shneiderman2020human] [Cai02012025, DBLP:journals/ijhci/BachKHBS24, 10.1145/3696449] [DBLP:journals/aisthci/RobertBMS20] [10.1145/3419764]
IS [DBLP:journals/jsis/LiHV08, DBLP:journals/ejis/SollnerHL16, DBLP:journals/tem/ThatcherMBAR11, DBLP:conf/hicss/ChopraW03, DBLP:journals/jcis/HoehleHG12, DBLP:journals/ejis/GrimsleyM07, Zhang2018, PARK2014153, DBLP:journals/isj/AllenCFK00, DBLP:journals/db/LiHV06, DBLP:journals/iam/MaoLD08, Tam2020, Tansley2007] None [acc470ea-2310-31ca-a502-a1b6f480aafd, DBLP:journals/jmis/BenbasatGP08] None
IS-AI [Saffarizadeh02072024, DBLP:journals/electronicmarkets/ThiebesLS21, DBLP:journals/jds/SchmidtBT20, Dennis03042023, https://doi.org/10.1111/jbl.12301, Das01112024, Troshani03092021, Korzyński12092024, DBLP:journals/giq/JanssenBEBJ20, DBLP:conf/wirtschaftsinformatik/TomitzaSSW23, info:doi/10.2196/15154] [DBLP:journals/csur/KaurURD23, DBLP:conf/hicss/LockeyGHS21, doi:10.5465/annals.2018.0057, DBLP:journals/electronicmarkets/YangW22] [DBLP:journals/electronicmarkets/LukyanenkoMS22] [DBLP:journals/isci/VirvouTT24]
SE [DBLP:journals/ese/BratthallJ02, DBLP:conf/esem/JalaliGS10, DBLP:conf/sigsoft/HassanLRGC00TOL24, DBLP:conf/icse/JohnsonBFFZ23, DBLP:journals/tse/LimamB10, DBLP:journals/software/BiancoLMT11, DBLP:journals/tse/CalinescuWGIHK18, DBLP:journals/jss/KeivanlooR14, DBLP:journals/ese/WangR16, DBLP:conf/icse/TrainerAR11, DBLP:conf/esem/BertramH0WS20, DBLP:journals/sigsoft/BeckerHPBKPDLRWGMSHMW06] [DBLP:journals/ese/HouJ23, DBLP:conf/sac/BuhnovaHIB23] None None
SE-AI [DBLP:journals/spe/AkbarKMRD24, DBLP:conf/icse/JohnsonBFFZ23] [DBLP:journals/csur/LiuLZJWS25, DBLP:journals/csur/ZhangDZ24, DBLP:journals/csur/KaurURD23, DBLP:journals/csur/LiQLDLPYZ23, DBLP:conf/hicss/LockeyGHS21] [DBLP:journals/electronicmarkets/LukyanenkoMS22] None
SE-CR [DBLP:conf/icse/NollerS0R22, DBLP:journals/tse/AlamiPCW22, DBLP:journals/tse/WinterBCHHNW23, DBLP:journals/pacmse/KouCW0024, DBLP:conf/chi/BoggustHSS22] None None None
@@ III-B1 Information Systems and Management (IS)
In IS research, trust has been studied since the 1990s [sollner2013we]. The articles we included in our review of IS literature were published in the time span from 2000 to 2024. Of the 24 research papers we analyzed, only four did not embed their studies in existing trust models. It was common for IS papers to customize existing trust models (13 out of 25 articles).
@@
Several articles discussed that trust is a dynamic concept that develops over time [DBLP:journals/jsis/LiHV08]. The factors and processes that lead to trust formation change gradually [Mayer1995], and can include aspects such as people’s perceptions about a company or brand [DBLP:journals/ejis/SollnerHL16]. This means that trust in a company can transfer to trust in a particular tool, a concept called trust transference [DBLP:journals/jmis/Stewart06, DBLP:journals/dss/WangSS13]. [Saffarizadeh02072024] applied this concept to AI agents and noted that trust in an AI creator can be transferred to its AI agent when there is a perceived meaningful association between them [Saffarizadeh02072024].
@@
IS articles clearly distinguish initial trust, which is important to create a trustworthy first impression of a new system and encourage adoption [DBLP:journals/jsis/LiHV08], from the before-mentioned process of trust formation and later stages such as knowledge-based trust, where the individual knows the other party well enough to predict the party’s behavior in a situation [DBLP:journals/electronicmarkets/ThiebesLS21]. From a tool perspective, the later phases of trust are sometimes referred to as postadoption, where aspects such as predictability and helpfulness of a tool [1985-30794-001] play an important role [DBLP:journals/tem/ThatcherMBAR11]. [DBLP:journals/jsis/McKnightCK02]’s model is frequently referred to as a common IS model for initial trust, whereas [DBLP:journals/misq/GefenKS03]’s model focuses on trust formation.
@@
In [DBLP:journals/jsis/McKnightCK02]’s model, based on [Mayer1995], trust in vendors of e-commerce websites is a multidimensional construct centered around trusting beliefs, that is, perceptions of the competence, benevolence, and integrity of the vendor and trusting intentions, that is, the willingness to depend or, in other words, the decision to make oneself vulnerable to the vendor [DBLP:journals/jsis/McKnightCK02]. In addition to a trust model, [DBLP:journals/jsis/McKnightCK02] also introduced instruments to measure the concepts in their model (see Figure 1 and Section IV).
@@
[DBLP:journals/misq/GefenKS03] models consumers’ online purchase intentions as a combination of perceived usefulness, ease-of-use, and trust in the vendor [DBLP:journals/misq/GefenKS03]. They show that online trust is built through (1) a belief that the vendor has nothing to gain by cheating, (2) a belief that there are safety mechanisms built into the web site, and (3) by having a typical interface, (4) one that is, moreover, easy to use.
@@
An important observation by [Mayer1995] is repeated in IS articles: humans differ in their general tendency to be willing to trust others [DBLP:journals/jsis/LiHV08], that is, their propensity [Mayer1995, DBLP:conf/hicss/ChopraW03] or disposition to trust [DBLP:journals/db/LiHV06]. These individual differences are a potential confounding factor in studies with developers investigating trust in AI assistants.
@@
Two other central constructs are trusting intention, which is the trustor’s willingness to depend on the trustee [DBLP:journals/jsis/LiHV08] and a person’s trusting beliefs [DBLP:journals/electronicmarkets/ThiebesLS21], that is favorable object-specific beliefs [DBLP:journals/tem/ThatcherMBAR11] that capture the trustor’s perception of whether the trustee has the desired attributes to be trusted [DBLP:journals/jsis/LiHV08] (see also [DBLP:journals/jsis/McKnightCK02]’s model). Moreover, IS literature discusses how the trusting beliefs proposed by [Mayer1995] (benevolence, ability, and integrity) can foster technology acceptance [DBLP:journals/jais/BenbasatW05]. [Troshani03092021] categorize those trusting beliefs into cognitive trust and affective trust [Troshani03092021] (see also [Komiak2006] [Komiak2006]). Cognitive trust captures a customer’s confidence or willingness to rely on a service provider’s competence and reliability while affective trust is the confidence that one places in another party on the basis of feelings generated by the level of care and concern as demonstrated by the other party. In summary, competence (i.e., ability in [Mayer1995]’s model) and predictability influence cognitive trust while benevolence and integrity influence affective trust [Troshani03092021].
@@
The IS literature discusses the validity of applying models of human trust to technology [DBLP:journals/jsis/LiHV08, DBLP:conf/hicss/ChopraW03, Saffarizadeh02072024], which is referred to as plausibility of trust in philosophy (see Section II-C). An aspect to consider is that humans have long been known to anthropomorphize technology [DBLP:journals/ijmms/NassMFRD95]. To study trust relationships between humans and human-like tools, researchers have adapted the concept of interpersonal trust to technology, for example, to study trust in automation or autonomous systems (see the description of [Lee2004Trust]’s trust model in Section III-B2). However, researchers have also argued that such models do not necessarily apply to GenAI tools due to their nondeterminism [info:doi/10.2196/15154].
@@
IS articles further discuss that increasing automation and complexity leads to systems being perceived as opaque and sophisticated, hence trust becomes more and more important [Lee2004Trust]. This situation leads to uncertainty and a perceived risk of using an automation tool, which could hinder adoption [DBLP:journals/jmis/BrownPR04, DBLP:journals/tem/ThatcherMBAR11]. However, [DBLP:conf/hicss/ChopraW03] also write that trust can only arise when there exists a state of dependence between the trustor and trustee, and when acting on this dependence entails risk, i.e., the trustor possesses uncertainty about the outcomes and vulnerability to a potential loss if the outcomes are undesirable [DBLP:conf/hicss/ChopraW03]. This is certainly true for software engineers who use GenAI tools to develop software. The role of risk and vulnerability appears also in the psychology literature, in particular [Mayer1995] (see Figure 1), as well as in the philosophical debate summarized in Section II-B.
@@
In the context of AI for decision-making, both mistakenly denied trust and unfounded trust can occur [DBLP:journals/jds/SchmidtBT20]. [DBLP:journals/jds/SchmidtBT20] summarize these situations as follows: If humans increasingly leverage AI to inform, derive, and justify decisions, it also becomes important to quantify when, how, why, and under which conditions they tend to overly trust or mistrust those systems [DBLP:journals/jds/SchmidtBT20]. This view corresponds to the concepts of undertrust and overtrust as described by [Lee2004Trust] [Lee2004Trust].
@@
Regarding trust quantification, [DBLP:journals/jds/SchmidtBT20] note that trust is measured differently in different research fields [DBLP:journals/jds/SchmidtBT20]. However, they also mention that a common approach used across disciplines is to ask study subjects about their trusting beliefs or trusting intentions toward a certain entity, for which standardized survey instruments exist (cf. [gefen2003managing]). An exemplary operationalization of trust in decision-making based on observed behavior is calculating the probability that users follow a model’s predictions by utilizing usage data [DBLP:conf/chi/Poursabzi-Sangdeh21, DBLP:journals/corr/abs-1901-08558]. This resembles our motivating example of software developers accepting code suggestions. We discuss trust operationalization and measurement in more detail in Section IV.